Five grams of sugar are added to a 70-gram serving of a breakfast cereal that is 10% sugar. What is the percent concentration of sugar in the resulting mixture?
Answer by lwsshak3(11628)

Five grams of sugar are added to a 70-gram serving of a breakfast cereal that is 10% sugar. What is the percent concentration of sugar in the resulting mixture?
**
amt of sugar in 70-gram serving=10%*70=7 grams
total amt of sugar in mixture=5+7=12 grams
total weight of mixture=70+5=75 grams
percent concentration of sugar in the resulting mixture=12/75=16%